# From Challenges and Pitfalls to Recommendations and Opportunities: Implementing Federated Learning in Healthcare

### v1

@@ 5.6 Fairness and Incentive
In FL for healthcare, research on fairness and incentive mechanisms is relatively underexplored, with only 53 delved into the intricacies of both fairness and incentive in FL for healthcare.
Fairness in FL refers to the equitable distribution of model performance across participants. In healthcare, it often means ensuring that ML models perform consistently across different healthcare providers and demographic groups or patient attributes. These fairness considerations are essential because disparities in model accuracy can lead to unequal treatment outcomes. 53 introduced several types of fairness, including horizontal fairness, where different hospitals receive comparable model accuracy, and vertical fairness, which focuses on ensuring that model performance is balanced across different demographic or medical attributes. They also proposed multilevel fairness, which seeks to address both client-level and attribute-level fairness simultaneously, and agnostic distribution fairness, aiming to generalize the model’s fairness to non-participating entities, such as hospitals outside the federation.
@@ 5.7.1 Gaps in Evaluation Metrics
The evaluation of FL models in healthcare heavily focuses on conventional ML metrics, such as accuracy, precision, AUC, sensitivity/recall, specificity, F1-score, Dice score, IoU, HD, and loss value 186. FL models are typically compared against classical centralized models or localized models, with ablation studies commonly used to isolate the impact of specific modifications. Most studies overlooked critical aspects unique to FL, such as communication overhead, resource consumption, privacy, and security concerns, thus failing to capture the complexity of FL systems.
FL involves frequent communication, which can introduce delays and increase costs. However, only a minority of studies (18%) included communication efficiency in their evaluation. Metrics such as communication cost, number of communication rounds, and latency are crucial for understanding the effectiveness of FL systems 109, 34.
Resource consumption is another crucial factor that has been underexplored, with only 12% of the reviewed studies measuring computational costs in FL evaluation. Key metrics such as training time 97, encryption time 187, CPU and memory consumption 57 are necessary for evaluating the computational efficiency of FL systems 47.
Privacy and security evaluations are fundamental for FL in healthcare. Despite this, only 16% of the reviewed studies assessed these aspects, with methods focusing on vulnerabilities to attacks such as model inversion attacks, and DP guarantees. For instance, 42 evaluated the influence of model inversion attacks on synthetic medical images in FL settings. 32 offered theoretical guarantees for DP to assess the system’s privacy resilience. 65 assessed various inversion attacks on medical images to measure and visualize potential data leakage in FL.
Scalability, a critical factor for the widespread adoption of FL in healthcare, has similarly been insufficiently evaluated. Variations in client numbers and patient populations need to be thoroughly assessed to ensure that FL systems can scale across large healthcare networks. Studies by 36 and 78 have demonstrated how FL models can handle varying clients and patient populations.
@@ 5.7.2 Insufficient Benchmarking
Only one study has reported standardized benchmarking, but it included a limited set of healthcare datasets and failed to integrate key constraints of FL in healthcare, particularly privacy, efficiency, and generalizability 52.
@@ 5.7.3 Lack of Interpretability
FL models allow decentralized data processing, but their black-box nature makes it difficult to understand how decisions are made. This is particularly critical in healthcare, where incorrect decisions, such as misdiagnoses or inappropriate treatment recommendations, can have severe consequences 22. The opacity of these models raises concerns about trust and accountability, as medical professionals must be able to explain and justify the decisions made by such systems 21.
The primary challenge in achieving interpretability in FL stems from the decentralized nature of data. Since each client’s data remains private, the global model lacks direct access to local datasets, making it harder to detect biases, noisy features, or irrelevant data points. Additionally, resource constraints (e.g., limited computational power and communication bandwidth) can hinder the integration of advanced interpretability techniques. Moreover, multiple stakeholders are involved in the decision-making process in FL: the central server needs to understand the significance of certain features, while clients must comprehend how their data contributes to the model’s performance.
Interpretable feature selection is an essential component of addressing these interpretability challenges in FL. By identifying the most relevant features and filtering out noisy or redundant data, FL models can not only improve performance but also increase transparency in their decision-making processes. In healthcare, this is particularly important, as clinicians need to understand which clinical factors the model considers most relevant. For instance, 62 utilized SHAP values to analyze the correlation between 20 clinical features and COVID-19 outcomes in FL settings, finding that eosinophil count had the greatest influence on predictions. 188 introduced a mutual information-based approach to select relevant features in a decentralized manner, while 189 proposed an unsupervised technique to detect outlier features and group similar ones via hierarchical clustering in FL.
In addition to feature selection, model-specific techniques such as tree-based FL provide further opportunities for interpretability by allowing models to function as “white-boxes” 135, 190. These methods leverage the model’s internal structure to explain its behavior. However, such approaches are often highly specific to particular types of data and may not be broadly applicable across different healthcare domains. Similarly, while gradient-based explanations and attention mechanisms offer another interpretability route, their effectiveness is sometimes limited due to weak correlations between these methods and the actual decision-making processes of the model 51, 89.
@@
  • Collaboration Between AI and Healthcare Professionals: Foster collaborations between AI experts and healthcare professionals to ensure that the interpretability techniques developed for FL models align with the needs and expectations of medical practitioners. Regular feedback loops could be established to fine-tune models for both accuracy and interpretability.
@@ 5.7.4 Poor Documentation and Reproducibility
  • Code & Model Release: Prioritize the release of well-documented code and trained models to facilitate independent performance evaluations and advance the field collaboratively.
  • Checklist: Create a comprehensive FL methodology checklist could improve documentation practices in future studies.
  • Piplines Documentation: Implement full-stack FL pipelines with documentation can simplify AI development for healthcare institutions, making advanced methods more accessible to users with information technology expertise.
  • Evaluation on Public Datasets: Encourage the use of public datasets for evaluation is also crucial to ensure fair comparisons and enhance the generalizability of results.
  • Framework: Extend existing FL frameworks rather than developing new ones should be promoted to reduce the risk of errors and support community-driven validation and improvement.

### v2

@@ 5.6 Fairness and Incentive
In FL for healthcare, research on fairness and incentive mechanisms is relatively underexplored, with only 55 delved into the intricacies of both fairness and incentive in FL for healthcare.
Fairness in FL refers to the equitable distribution of model performance across participants. In healthcare, it often means ensuring that ML models perform consistently across different healthcare providers and demographic groups or patient attributes. These fairness considerations are essential because disparities in model accuracy can lead to unequal treatment outcomes. 55 introduced several types of fairness, including horizontal fairness, where different hospitals receive comparable model accuracy, and vertical fairness, which focuses on ensuring that model performance is balanced across different demographic or medical attributes. They also proposed multilevel fairness, which seeks to address both client-level and attribute-level fairness simultaneously, and agnostic distribution fairness, aiming to generalize the model’s fairness to non-participating entities, such as hospitals outside the federation.
@@ 5.7.1 Gaps in Evaluation Metrics
The evaluation of FL models in healthcare heavily focuses on conventional ML metrics, such as accuracy, precision, AUC, sensitivity/recall, specificity, F1-score, Dice score, IoU, HD, and loss value 196. FL models are typically compared against classical centralized models or localized models, with ablation studies commonly used to isolate the impact of specific modifications. Most studies overlooked critical aspects unique to FL, such as communication overhead, resource consumption, privacy, and security concerns, thus failing to capture the complexity of FL systems.
FL involves frequent communication, which can introduce delays and increase costs. However, only a minority of studies (18%) included communication efficiency in their evaluation. Metrics such as communication cost, number of communication rounds, and latency are crucial for understanding the effectiveness of FL systems 110, 36.
Resource consumption is another crucial factor that has been underexplored, with only 12% of the reviewed studies measuring computational costs in FL evaluation. Key metrics such as training time 98, encryption time 197, CPU and memory consumption 29 are necessary for evaluating the computational efficiency and understanding the feasibility of FL systems 49. Without these metrics, it is challenging to comprehensively evaluate the trade-offs between performance and resource requirements, limiting the ability to assess the feasibility and practicality of deploying FL systems in resource-constrained healthcare environments.
Privacy and security evaluations are fundamental for FL in healthcare. Despite this, only 16% of the reviewed studies assessed these aspects, with methods focusing on vulnerabilities to attacks such as model inversion attacks, and DP guarantees. For instance, 44 evaluated the influence of model inversion attacks on synthetic medical images in FL settings. 34 offered theoretical guarantees for DP to assess the system’s privacy resilience. 66 assessed various inversion attacks on medical images to measure and visualize potential data leakage in FL.
Scalability, a critical factor for the widespread adoption of FL in healthcare, has similarly been insufficiently evaluated. Variations in client numbers and patient populations need to be thoroughly assessed to ensure that FL systems can scale across large healthcare networks. Studies by 38 and 79 have demonstrated how FL models can handle varying clients and patient populations.
@@ 5.7.2 Insufficient Benchmarking
Only one study has reported standardized benchmarking, but it included a limited set of healthcare datasets and failed to integrate key constraints of FL in healthcare, particularly privacy, efficiency, and generalizability 54.
@@ 5.7.3 Lack of Interpretability
FL models allow decentralized data processing, but their black-box nature makes it difficult to understand how decisions are made. The opacity of these models raises concerns about trust and accountability, as medical professionals must be able to explain and justify the decisions made by such systems 22, 21.
The primary challenge in achieving interpretability in FL stems from the decentralized nature of data. Since each client’s data remains private, the global model lacks direct access to local datasets, making it harder to detect biases, noisy features, or irrelevant data points. Additionally, privacy-preserving mechanisms, such as DP, can obscure data details, further complicating efforts to generate meaningful explanations. Moreover, multiple stakeholders are involved in the decision-making process in FL: the central server needs to understand the significance of certain features to ensure reliable global updates, while clients must comprehend how their data contributes to the model’s performance. These multi-level interpretability needs, combined with resource constraints (e.g., limited computational power and communication bandwidth), present unique challenges for integrating advanced interpretability techniques in FL.
Interpretable feature selection is an essential component of addressing these interpretability challenges in FL. By identifying the most relevant features and filtering out noisy or redundant data, FL models can not only improve performance but also increase transparency in their decision-making processes. In healthcare, this is particularly important, as clinicians need to understand which clinical factors the model considers most relevant. For instance, 63 utilized SHAP values to analyze the correlation between 20 clinical features and COVID-19 outcomes in FL settings, finding that eosinophil count had the greatest influence on predictions. 198 introduced a mutual information-based approach to select relevant features in a decentralized manner, while 199 proposed an unsupervised technique to detect outlier features and group similar ones via hierarchical clustering in FL.
In addition to feature selection, model-specific techniques such as tree-based FL provide further opportunities for interpretability by allowing models to function as “white-boxes” 137, 200. These methods leverage the model’s internal structure to explain its behavior. However, such approaches are often highly specific to particular types of data and may not be broadly applicable across different healthcare domains. Similarly, while gradient-based explanations and attention mechanisms offer another interpretability route, their effectiveness is sometimes limited due to weak correlations between these methods and the actual decision-making processes of the model 53, 90.
@@
  • Privacy-Preserving Explainability: Develop interpretability techniques that align with privacy-preserving requirements in FL, such as DP-aware SHAP values or SMPC for feature importance analysis. These methods should balance transparency with the need to protect sensitive data.
@@ 5.7.4 Poor Documentation and Reproducibility
  • Code & Model Release: Prioritize the release of well-documented code and trained models to facilitate independent performance evaluations and advance the field collaboratively. However, to balance transparency with privacy, ensure that shared models incorporate privacy-preserving techniques such as DP or HE to safeguard sensitive information.
  • Checklist: Create a comprehensive FL methodology checklist to improve documentation practices in future studies. The checklist should include guidelines for maintaining transparency while adhering to data privacy standards.
  • Pipelines Documentation: Implement full-stack FL pipelines with documentation to simplify AI development for healthcare institutions, making advanced methods more accessible to users with information technology expertise. Modular and privacy-aware pipeline designs are recommended to reduce the risk of exposing sensitive details.
  • Evaluation on Public Datasets: Encourage the use of public datasets for evaluation to ensure fair comparisons and enhance the generalizability of results. Anonymized or synthetic datasets could also be employed when real-world data cannot be shared openly due to privacy concerns.
  • Framework: Extend existing FL frameworks rather than developing new ones to reduce the risk of errors and support community-driven validation and improvement. Incorporating privacy-preserving modules into these frameworks can address both security and transparency needs effectively.
  • Privacy & Transparency Balance: Emphasize the importance of balancing open-sourcing efforts with robust privacy measures. Transparency initiatives should focus on sharing aggregated insights, generalizable methodologies, and anonymized results while safeguarding patient data integrity.